# Avoiding Catastrophe in Online Learning by Asking for Help

### v4

@@ B.2 Main proof
(The $+1$ is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Therefore
$\operatorname*{\mathbb{E}}[|Q_{T}|]=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{X}|]$
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{X}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+\sum_{y\in\mathcal{Y}}O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
@@
Theorem 5.1 follows from Lemmas B.4 and B.7:
@@
We then perform some arithmetic to get Theorem 5.2:
@@
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{dL}{\sigma}T^{1-\frac{2n}{2n+1}-\frac{2}{2n+1}}\log(1/\varepsilon)\log T\right)$
$=$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
and
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\log(T^{\frac{2n}{2n+1}})\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\boldsymbol{x})^{n}\right)$
$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\boldsymbol{x})^{n}\right)$
$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\Big{)}\right)$
If we instead use the second bound from Lemma B.4, the same arithmetic gives us:
See 5.3
@@ B.3 Adaptive adversaries
In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of Slivkins et al. (2019) or Chapter 21 of Shalev-Shwartz & Ben-David (2014)). However, Russo et al. (2024) state their result only for oblivious adversaries. In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in Russo et al. (2024)) must also hold for adaptive adversaries. In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well. For this rest of Section B.3, lemma numbers refer to the numbering in Russo et al. (2024).
@@
Recall from Section B.1 that Russo et al. (2024) allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity. Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2. More precisely, on each time step $t$, the algorithm samples a Bernoulli random variable $X_{t}\sim\text{Ber}(p)$ and queries if $X_{t}=1$. The key idea is that $X_{t}$ is independent of events on previous time steps. Thus even conditioning on the history up to time $t$, for any for any random variable $Y_{t}$ we can write
@@
$\operatorname*{\mathbb{E}}[\ell_{t}(i_{t})]=(1-p)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+p\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]=\left(1-\frac{k}{T}\right)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+\frac{k}{T}\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]$
@@
The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by ^$\hat{\ell}_{t}=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$. To do so, we must argue that standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries. Specifically, both algorithms must have the same weight vectors on every time step, and the only difference should be that HedgeWithQueries takes the optimal action on each time step independently with probability $p$ (and otherwise behaves the same as standard Hedge). On time steps with $X_{t}=0$, standard Hedge observes ^$\hat{\ell}_{t}(i)=0$ for all actions $i$ and thus makes no updates, and HedgeWithQueries makes no updates by definition. On time steps with $X_{t}=1$, both algorithms perform the typical updates $w_{t+1}(i)=w_{t}(i)\cdot\exp(-\eta(\hat{\ell}_{t}(i)-\hat{\ell}_{t}(i_{t}^{*})))$. Thus the weight vectors are the same for both algorithms on every time step. Furthermore, HedgeWithQueries takes the optimal action at time $t$ iff $X_{t}=1$, which occurs independently with probability $p$ on each time step. Thus standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries.
@@
We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of Shalev-Shwartz & Ben-David (2014)). For each action $y$, we will learn a binary classifier which predicts whether action $y$ is the mentor’s action. Formally, for each $y\in\mathcal{Y}$, define the policy class $\Pi_{y}=\{\pi_{y}:\pi\in\Pi\text{ and }\pi_{y}(x)=\mathbf{1}(\pi(x)=y))\ \forall x\in\mathcal{X}\}$. In words, for each policy $\pi:\mathcal{X}\to\mathcal{Y}$ in $\Pi$, there exists a policy $\pi_{y}:\mathcal{X}\to\{0,1\}$ in $\Pi_{y}$ such that $\pi_{y}(x)=\mathbf{1}(\pi(x)=y)$ for all $x\in\mathcal{X}$.
Algorithm 3 runs one copy of our binary-action algorithm Algorithm 1 for each action $y\in\mathcal{Y}$. At each time step $t$, the copy for action $y$ returns an action $b_{t}^{y}$, with $b_{t}^{y}=1$ indicating a belief that $y=\pi^{m}(x_{t})$ and $b_{t}^{y}=0$ indicating a belief that $y\neq\pi^{m}(x_{t})$. (Note that $b_{t}^{y}=\tilde{y}$ is also possible, indicating that the mentor was queried.)
@@
1:function AvoidCatastropheManyActions($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
2:     for $y\in\mathcal{Y}$ do
3:         if $\Pi$ has VC dimension $d$ then
4:              $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
5:         else if $\Pi$ has Littlestone dimension $d$ then
6:              $\tilde{\Pi}_{y}\leftarrow$ any adversarial $\varepsilon$-cover of size at most $(eT/d)^{d}$
7:     for $t$ from $1$ to $T$ do
8:         for $y\in\mathcal{Y}$ do
9:              $b_{t}^{y}\leftarrow$ action from running one step of Algorithm 1 on $\Pi_{y}$ (with the same $T,\varepsilon,d$)
10:         if $b_{t}^{y}\neq\tilde{y}$ $\forall y\in\mathcal{Y}$ and $\exists a\in\mathcal{Y}:b_{t}^{y}=1$ then
11:              Take any action $y$ with $b_{t}^{y}=1$
12:         else
13:              Query the mentor
@@
Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$. Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\right)\right)$
@@
    Let $\pi_{t}^{y}$ and $X_{t}^{y}$ refer to the values of $\pi_{t}$ and $X_{t}$ for copy $y$ of Algorithm 1.
@@
    Let $M_{T}^{y}=\{t\in[T]:b_{t}^{y}\neq\pi^{my}(x_{t})\}$ be the set of time steps where $\pi_{t}^{y}$ does not correctly determine whether the mentor would take action $y$ and let $M_{T}=\{t\in[T]:y_{t}\neq\pi^{m}(x_{t})\}$ be the set of time steps where the agent’s action doesn’t match the mentor’s.
@@
The argument is similar to the proof of Lemma B.2. If $\mu_{t}^{m}(x_{t})\neq\mu_{t}(x_{t},y_{t})$, then $y_{t}=y$ for some $y\in\mathcal{Y}$ where $b_{t}^{y}=1$. Therefore copy $y$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{y}(x_{t})=1$. Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in X_{t}^{y}:\pi_{t}^{y}(x_{t})=y}||x_{t}-x||$. Then $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$ and $y^{\prime}=\pi_{t}^{y}(x_{t})=1$.
By construction of $X_{t}^{y}$, $y^{\prime}=\pi^{my}(x^{\prime})$ so $\pi^{my}(x^{\prime})=1$ which implies $\pi^{m}(x^{\prime})=y$. Then by the local generalization assumption,
@@
See C.1
Proof.
Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\big{)}\big{)}$ queries in expectation, so by linearity of expectation, the expected number of queries made by Algorithm 3 is $O\big{(}|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\big{)}\big{)}$. Using the same argument as in the proof of Lemma B.4 (with Lemma C.2 replacing Lemma B.2), we get
$R_{T}=$ $\ \prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
$\leq$ $\ |M_{T}|L\varepsilon^{1/n}$
Then by Lemma C.1, $R_{T}\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}|M_{T}^{y}|$. Taking the expectation and applying Lemma B.1 to each $M_{T}^{y}$ gives us
$\operatorname*{\mathbb{E}}[R_{T}]\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|\mathcal{Y}|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
Plugging in $\varepsilon=T^{\frac{-2n}{2n+1}}$ produces the desired bounds (see the arithmetic in the proof of Theorem 5.2 in Appendix B). ∎
One can use the same argument to obtain a multi-action additive regret bound: simply modify the proof of Theorem C.1 to use $\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq|M_{T}|L\varepsilon^{1/n}$ instead of $R_{T}\leq|M_{T}|L\varepsilon^{1/n}$.
Corollary C.1.1.
Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$. Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$

### v5

@@ B.2 Main proof
(The $+1$ is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Since $\hat{X}_{2}\subseteq\{x_{1},\dots,x_{T}\}\subseteq\cup_{y\in\mathcal{Y}}X_{y}$, we have $|\hat{X}_{2}|\leq\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|$ by the union bound. Therefore
$\operatorname*{\mathbb{E}}[|Q_{T}|]\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}|]$
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{Q}_{2}|]$
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}[|\hat{X}_{2}|]$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+\operatorname*{\mathbb{E}}[|\hat{Q}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
[…]
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+1+O\left(\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}+1\right)$
$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(T+1/\varepsilon)+\frac{\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]}{\varepsilon}\right)$
@@
Theorem 5.1 follows from Lemmas B.3 and B.6:
@@
We then perform some arithmetic to obtain Theorem 5.2:
@@
$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T^{1-\frac{2n}{2n+1}-\frac{2}{2n+1}}\left(\log T+\log(T^{\frac{2n}{2n+1}})\right)\right)$
$=$ $\ O\left(\frac{dL}{\sigma\mu_{0}^{m}}T^{\frac{-1}{2n+1}}\log T\right)$
and similarly for $\operatorname*{\mathbb{E}}[R_{T}^{+}]$. For $\operatorname*{\mathbb{E}}[|Q_{T}|]$,
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\left(\log T+\log(T^{\frac{2n}{2n+1}})\right)+T^{\frac{2n}{2n+1}}\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\right)$
$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\right)$
$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\Big{)}\right)$
@@ B.3 Adaptive adversaries
In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of Slivkins et al. (2019) or Chapter 21 of Shalev-Shwartz & Ben-David (2014)). However, Russo et al. (2024) state their result only for oblivious adversaries. In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in Russo et al., 2024) must also hold for adaptive adversaries. In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well. For this rest of Section B.3, lemma numbers refer to the numbering in Russo et al. (2024).
@@
Recall from Section B.1 that Russo et al. (2024) allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity. Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2. More precisely, on each time step $t$, the algorithm samples $X_{t}\sim\text{Bernoulli}(p)$ and queries if $X_{t}=1$. The key idea is that $X_{t}$ is independent of events on previous time steps. Thus even conditioning on the history up to time $t$, for any for any random variable $Y_{t}$ we can write
@@
$\operatorname*{\mathbb{E}}[\ell_{t}(i_{t})]=$ $\ (1-p)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+p\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]$
$=$ $\ \left(1-\frac{k}{T}\right)\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{0})]+\frac{k}{T}\operatorname*{\mathbb{E}}[\ell_{t}(i_{t}^{*})]$
@@
The other tricky part of the proof is applying Observation 3.3 using a new loss function $\hat{\ell}$ defined by ^$\hat{\ell}_{t}(i)=\frac{T}{\hat{k}}(\ell_{t}(i)-\ell_{t}(i_{t}^{*}))\mathbf{1}(X_{t}=1)$. To do so, we must argue that standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries. Specifically, both algorithms must have the same weight vectors on every time step, and the only difference should be that HedgeWithQueries takes the optimal action on each time step independently with probability $p$ (and otherwise behaves the same as standard Hedge). On time steps with $X_{t}=0$, standard Hedge observes ^$\hat{\ell}_{t}(i)=0$ for all actions $i$ and thus makes no updates, and HedgeWithQueries makes no updates by definition. On time steps with $X_{t}=1$, both algorithms perform the typical updates $w_{t+1}(i)=w_{t}(i)\cdot\exp(-\eta(\hat{\ell}_{t}(i)-\hat{\ell}_{t}(i_{t}^{*})))$. Thus the weight vectors are the same for both algorithms on every time step. Furthermore, HedgeWithQueries takes the optimal action at time $t$ iff $X_{t}=1$, which occurs independently with probability $p$ on each time step. Thus standard Hedge run on $\hat{\ell}$ is the “counterpart without queries” of HedgeWithQueries. Note that since $\hat{\ell}$ is itself a random variable, the law of iterated expectation is necessary to formalize this.
@@
We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of Shalev-Shwartz & Ben-David, 2014). For each action $y$, we will learn a binary classifier which predicts whether action $y$ is the mentor’s action. Formally, for each $y\in\mathcal{Y}$, define the policy class $\Pi_{y}=\{\pi_{y}:\pi\in\Pi\text{ and }\pi_{y}(x)=\mathbf{1}(\pi(x)=y))\ \forall x\in\mathcal{X}\}$. In words, for each policy $\pi:\mathcal{X}\to\mathcal{Y}$ in $\Pi$, there exists a policy $\pi_{y}:\mathcal{X}\to\{0,1\}$ in $\Pi_{y}$ such that $\pi_{y}(x)=\mathbf{1}(\pi(x)=y)$ for all $x\in\mathcal{X}$.
Algorithm 3 runs one copy of our binary-action algorithm (Algorithm 1) for each action $y\in\mathcal{Y}$. At each time step $t$, the copy for action $y$ returns an action $b_{t}^{y}$, with $b_{t}^{y}=1$ indicating a belief that $y=\pi^{m}(x_{t})$ and $b_{t}^{y}=0$ indicating a belief that $y\neq\pi^{m}(x_{t})$. (Note that $b_{t}^{y}=\tilde{y}$ is also possible, indicating that the mentor was queried.)
@@
  Inputs: $T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$
  for $y\in\mathcal{Y}$ do
     if $\Pi_{y}$ has VC dimension $d$ then
        $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
     else if $\Pi$ has Littlestone dimension $d$ then
        $\tilde{\Pi}_{y}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$
  for $t$ from $1$ to $T$ do
     for $y\in\mathcal{Y}$ do
        $b_{t}^{y}\leftarrow$ action at time $t$ from the copy of Algorithm 1 running on $\Pi_{y}$ (with the same $T,\varepsilon,d$)
     if $b_{t}^{y}\neq\tilde{y}$ $\forall y\in\mathcal{Y}$ and $\exists a\in\mathcal{Y}:b_{t}^{y}=1$ then
        Take any action $y$ with $b_{t}^{y}=1$
     else
        Take an arbitrary action in $\mathcal{Y}$
@@
Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$. Then for any $T\in\mathbb{N}$, Algorithm 3 with $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}^{\times}\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma\mu_{0}^{m}}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}\left[R_{T}^{+}\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\right)\right)$
@@
    Recall that $S_{t}$ refers to the value of $S$ at the start of time step $t$ in Algorithm 1. Let $\pi_{t}^{y}$ and $S_{t}^{y}$ refer to the values of $\pi_{t}$ and $S_{t}$ for copy $y$ of Algorithm 1.
@@
    Let $M_{T}^{y}=\{t\in[T]:b_{t}^{y}\neq\pi^{my}(x_{t})\}$ be the set of time steps where $\pi_{t}^{y}$ does not correctly determine whether the mentor would take action $y$ and let $M_{T}=\{t\in[T]:y_{t}\neq\pi^{m}(x_{t})\}$ be the set of time steps where the agent’s action does not match the mentor’s.
@@
The argument is similar to the proof of Lemma B.2. If $\mu_{t}^{m}(x_{t})\neq\mu_{t}(x_{t},y_{t})$, then $y_{t}=y$ for some $y\in\mathcal{Y}$ where $b_{t}^{y}=1$. Therefore copy $y$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{y}(x_{t})=1$. Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in S_{t}^{y}:\pi_{t}^{y}(x_{t})=y}||x_{t}-x||$. Then $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$ and $y^{\prime}=\pi_{t}^{y}(x_{t})=1$.
By construction of $S_{t}^{y}$, $y^{\prime}=\pi^{my}(x^{\prime})$ so $\pi^{my}(x^{\prime})=1$ which implies $\pi^{m}(x^{\prime})=y$. Then by the local generalization assumption,
@@
We now proceed to the proof of Theorem C.1.
Proof of Theorem C.1.
Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\big{)}\big{)}$ queries in expectation. Thus by linearity of expectation the expected number of queries made by Algorithm 3 is $O\big{(}|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{\mathbb{E}}[\operatorname*{diam}(\boldsymbol{x})^{n}]\big{)}\big{)}$. [^13] This is an overestimate because the agent makes at most one query per time step, even if multiple copies request a query. Similar to the proof of Lemma B.3, we have
[…]
Since each copy satisfies the conditions of Lemma B.1, we get
$\operatorname*{\mathbb{E}}[R_{T}^{+}]\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|\mathcal{Y}|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
Since $\lim_{T\to\infty}\varepsilon=0$, there exists $T_{0}$ such that $L\varepsilon^{1/n}\leq\mu_{0}^{m}/2$ for all $T\geq T_{0}$. Then by Lemma A.1,
$\operatorname*{\mathbb{E}}[R_{T}^{\times}]\leq\frac{\operatorname*{\mathbb{E}}[R_{T}^{+}]}{\mu_{0}^{m}/2}\in O\left(|\mathcal{Y}|L\varepsilon^{1/n}\frac{d}{\sigma\mu_{0}^{m}}T\varepsilon\log(1/\varepsilon)\log T\right)$ (3)
Plugging $\varepsilon=T^{\frac{-2n}{2n+1}}$ to the bounds above on $\operatorname*{\mathbb{E}}[R_{T}^{+}]$ and $\operatorname*{\mathbb{E}}[R_{T}^{\times}]$ yields the desired bounds (see the arithmetic in the proof of Theorem 5.2 in Appendix B). ∎